# Shavette: Safe Undervolting of Neural Network Accelerators via Algorithm-level Error Detection

### v1

@@ V-B Error Detection
Figure 3, depicts the errors appearing in computations, i.e., actual error rate, error rate from ABFT detections and overall prediction errors (accuracy loss) of the model. Since error propagate from one layer to another, comparing ABFT error detection rate against actual errors per each operations is misleading. Experiments on single operation, shows very high (close to 100%) computational detection rate for ABFT, similiar to results on in [safarpour2021high]. The output error was reported if the difference between the GPU accelerated outputs and the reference was over 1e-13. The error threshold can be made slightly tighter to ensure more reliability, however, that would result in false positives being detected constantly, even at stock voltage and frequency settings.The DNN accuracy was compared against baseline (not undervolted). Notice, the accuracy of model does not degrade until the computational error rate grows substantially. Due to inherent fault-tolerance of DNNs a few hundred errors at vicinity of PoFF does not impact the accuracy. Nonetheless, since ABFT detects the computational errors early enough, we do not need to rely on fault-tolerance feature of DNNs for reliability.
@@ V-C Overheads
The power consumption at all voltages with ABFT enabled is slightly lower (around 1%) than when it is disabled, meaning ABFT overheads manifest in inference time rather than power. That is likely due to ABFT related calculations imposing more idle periods on processing units. As Table  I shows, this translates into slightly longer inference time. Hence, we can get an overall energy overhead estimate by comparing the energy per inferences between ABFT enabled and disabled runs. Table II provides a summary of overheads in terms of inference time.
Naive integration of ABFT into the DNN’s model results in large overheads, however, with small tweaks and code optimization the overhead share becomes insignificant, e.g., $<3.9\%$ for VGG-16. As mentioned earlier, the overhead ratio of adding ABFT grows sub-linearly with respect to the model size, hence, for LeNet-5 we observe larger overheads compared to VGG-16 due to the smaller model size.
@@
Here, we mainly consider VGG-16 for analysis due to it larger model size compared to LeNet and it is more representative of models implemented in real world in data centers. Considering VGG-16 with default clock frequency and voltage (at 1780 MHz and 960 mV), the energy overhead of ABFT is as low as 1%. With same frequency but voltage set to 880 mV, the energy overhead is almost the same, i.e., 1.3%. Setting the frequency to the lowest, i.e., 1680 MHz, and at the lowest stable around 805 mV, the energy overhead of ABFT enabled is slightly higher, i.e., 1.9%. This is a small penalty for saving almost a quarter of the energy of each inference.
TABLE II: Time required in seconds for 100 inferences with LeNet-5 and VGG-16 at different clock frequencies
VGG-16 @ 1820 MHz VGG-16 @ 1780 MHz VGG-16 @ 1680 MHz LeNet-5
ABFT disabled 17.190 17.519 18.342 8.08
ABFT enabled 17.808 18.136 18.986 8.7
Time overhead 3.6% 3.5% 3.5% 7.8%
VI Related works
@@
In [gundi2020effort], a Tensor Processing Unit (TPU) architecture, named ”EFFORT”, is proposed. The TPU is equipped with timing error detection circuits within critical paths of Multiply and Accumulate (MAC) units to enable reduced voltage operations. EFFORT achieves high performance and energy efficiency by employing TED systems and in-situ clock gating techniques to address. To minimize hardware overheads, not all critical paths can be equipped with TEDs and hence a 2% average accuracy drop across various DNN datasets was reported.
@@
Similar to EFFORT, in [fan2022design] a lightweight error detection circuit is introduced to detect timing errors from reduced voltage operations. However, circuit level techniques substantially complicate design and test process and are not applicable to already fabricated chips. Notice, that our proposed solution does not compromise the accuracy of DNN. MATIC [kim2018matic] combines algorithm-level and circuit-level techniques to enable efficient low-voltage operation of neural accelerators. No online error detection is utilized, instead, MATIC addresses the challenge of timing errors based on speculative fault model and error-aware training. Hence, once the fault model no longer is applicable, the accuracy can not be ensured.
One of the earliest works on algorithmic approaches for for enabling reduced voltage operation was introduced in [zamani2019greenmm]. However, they limit the application only to general matrix multiplication. We solved this limitation by developing ABFT enabled CNN models to demonstrate possibility for improved energy efficiency and reliable reduced voltage acceleration of CNNs.
VII Discussion and future works
We showcased the solution for two DNN models running on commodity GPU used as an AI accelerator. Although, we implemented the models from scratch, the necessary modifications can be made to the DNN automatically using development tools such as TensorFlow or Keras [zhao2020ft] and inserting the ABFT schemes into the models.
In a data center the Power Usage Effectiveness (PUE) is always above 1.0 and typically around 1.3 or even 1.8, i.e., 30% to 80% more power dissipated to support computing equipment. Removing voltage margins not only provides for direct power saving in the accelerators, but also leads to indirect cost savings coming from lower costs of power delivery network, lower demands of cooling and ultimately higher life time.
This work demonstrated possibility of utilizing ABFT for reducing power consumption of DNN inference. However, this work can be extended for training as well [zhao2020ft].
@@ VII Discussion and future works
Although the solution was only showcased on a GPU, we believe similar approaches can be applied to other DNN accelerators. Due to software based nature of this work, Shavette, is more cost efficient compared to hardware based ones, e.g., Razor [ernst2003razor]. Although, errors from control path appears after data-path (see Fig. 3) [safarpour2021high], the control path can be augmented by TED with low overheads while the data-path is protected by ABFT, to ensure reliability of all sub-components within the design.
VIII Conclusion
@@
This study demonstrates the existence of a significant voltage margin in the default settings of a commercial GPU and possibility of reliably removing it. In computations of DNN models the margin can be safely optimized by incorporating ABFT schemes. For the used VGG-16 The time and energy overheads from ABFT are low, e.g., around 3.5% and 1.1% respectively. Nearly a quarter of energy could be saved without sacrificing accuracy of the neural network. One of the issues that was verified in the context of this paper is that when the voltage is reduced the control path fails long after the data path. Nonetheless, the control path can be aggressively pipelined or be operated at slightly higher voltage such that it is ensured that data path always fails before control path. We believe the results indicate a promising avenue for enhancing the energy efficiency of AI accelerators, without costly circuit techniques, cell characterization or having to hack CAD tools.

### v2

@@ 4.2. Error Coverage
Figure 4, depicts the errors appearing in computations, i.e., actual error rate, error rate from ABFT detections and overall prediction errors (accuracy loss) of the model. Since errors propagate from one layer to another, comparing ABFT error detection rate against actual errors per each operations is misleading. Experiments on single operation, shows very high (close to 100%) computational detection rate for ABFT, similiar to results on in (Safarpour et al., 2021b). Errors were reported if difference between the GPU accelerated outputs and the reference was over 1e-13. The error threshold can be made slightly tighter to ensure more reliability, however, that would result in false positives being detected constantly, even at stock voltage and frequency settings.The DNN accuracy was compared against baseline (not undervolted). Notably, due to inherent fault-tolerance of DNNs a few hundred errors at vicinity of PoFF does not impact the accuracy, as it is observed in Fig. 4. Nonetheless, since ABFT detects the computational errors early enough, we do not need to rely on fault-tolerance feature of DNNs for ensuring reliability.
@@ 4.3. Overheads
The power consumption at all voltages with ABFT enabled is slightly lower (around 1%) than when it is disabled, meaning ABFT overheads manifest in inference time rather than power. That is likely due to ABFT related calculations imposing more idle periods on processing units. As Table  1 shows, this translates into slightly longer inference time. Overall energy overhead estimated by comparing the energy per inferences between ABFT enabled and disabled runs. Table 2 provides a summary of overheads in terms of inference time.
@@
Naive integration of ABFT into the DNN model results in large overheads, however, with small tweaks and code optimization the overhead share becomes insignificant, e.g., $<3.9\%$ for VGG-16. Considering VGG-16 with default clock frequency and voltage (at 1780 MHz and 960 mV), the energy overhead of ABFT is as low as 1%. With same frequency but voltage set to 880 mV, the energy overhead is almost the same, i.e., 1.3%. Setting the frequency to the lowest, i.e., 1680 MHz, and at the lowest stable around 805 mV, the energy overhead of ABFT enabled is slightly higher, i.e., 1.9%. This is a small penalty for saving almost a quarter of the energy of each inference.
Table 2. Inference times VGG-16 at different clock frequencies.
Configuration VGG-16 VGG-16 VGG-16
@ 1820 MHz @ 1780 MHz @ 1680 MHz
ABFT disabled ($ms)$ 171.90 175.19 183.42
ABFT enabled ($ms$) 178.08 181.36 189.86
Time overhead (%) $\approx 3.5\%$
5. Related works
@@
In (Gundi et al., 2020), a Tensor Processing Unit (TPU) architecture, named ”EFFORT”, is proposed. The TPU is equipped with timing error detection circuits within critical paths of Multiply and Accumulate (MAC) units to enable reduced voltage operations. EFFORT achieves high performance and energy efficiency by employing TED systems and in-situ clock gating techniques to address. To minimize hardware overheads, not all critical paths can be equipped with TEDs and hence a 2% average accuracy drop across various DNN datasets was reported.
@@
Similar to EFFORT, in (Fan et al., 2022) a lightweight error detection circuit is introduced to detect timing errors from reduced voltage operations. However, circuit level techniques substantially complicate design and test process and are not applicable to already fabricated chips. Notice, that our proposed solution does not compromise the accuracy of DNN. MATIC (Kim et al., 2018) combines algorithm-level and circuit-level techniques to enable efficient low-voltage operation of neural accelerators. No online error detection is utilized, instead, MATIC addresses the challenge of timing errors based on speculative fault model and error-aware training. Hence, once the fault model no longer is applicable, the accuracy can not be ensured.
@@
This work can be extended for training as well. The solution was only showcased on a GPU, while a similar approach can be applied to other DNN accelerators, e.g., Google TPUs. Due to software based nature of this work, Shavette, is more cost efficient compared to hardware based methods, e.g., Razor (Ernst et al., 2003).
6. Conclusion
@@
This study demonstrates the existence of a significant voltage margin in the default settings of a commercial GPU and possibility of reliably removing it. In computations of DNN models, the margin can be safely optimized by incorporating ABFT schemes. The time and energy overheads from ABFT are low, e.g., around 3.5% and 1.1% respectively. Nearly a quarter of energy could be saved without sacrificing accuracy of the neural network. The results indicate a promising avenue for enhancing the energy efficiency of AI accelerators, without expensive HW techniques or having to compromise reliability.